# On the Non-Orientable $3$- and $4$-Genera of a Knot: Connections and Comparisons

### v1

@@ 5. Ordinary Genus Gaps: Pretzel Knots
Proof.
The fact that $\gamma_{3}(P)=2n$ follows directly from the work of Ichihara and Mizushima [27, Theorem 1.2].
@@
As mentioned at the beginning of this section, despite the large gaps between $\gamma_{3}$ and $\gamma_{4}$ for the knots in Proposition 1.7, these examples do not exhibit similarly large gaps between $\hat{\gamma}^{\pm}_{3}$ and $\hat{\gamma}^{\pm}_{4}$.
@@
$\hat{\gamma}^{\pm}_{3}(P)-\hat{\gamma}^{\pm}_{4}(P)\leq 2.$
We conjecture that $\hat{\gamma}^{\pm}_{3}(P)-\hat{\gamma}^{\pm}_{4}(P)=0$. The conjecture would hold if the Möbius band filling constructed in the proof of Proposition 1.7 were the unique such band.
@@
Let $F_{A}$ (resp. $F_{B}$) be the all-$A$ (resp. all-$B$) state surface for $P$. Using Lemma 2.4, we compute that
$e(F_{A})=2k+\sum_{i=1}^{n-1}p_{i}\quad\text{and}\quad e(F_{B})=-4r-2-\sum_{i=1}^{n-1}p_{i}.$
A direct computation on the diagram yields
@@
6. Euler-Normalized Genus Gaps: Slice Pretzel Knots
In this section, we prove Proposition 1.8 by exhibiting an infinite family of non-alternating slice pretzel knots with negative Euler-normalized non-orientable $3$-genus 22. The techniques deployed in the proof are interesting more for their suggestion of what work would be necessary for the computation of the Euler-normalized genera of adequate knots and the proof of Conjecture 1.6 than for the result itself which, as noted above, would also follow from Theorems 1.3 and 1.4. The idea is reminiscent of the computation of the $3$-dimensional geography of the trefoil (Example 2.5) and of the $5_{2}$ knot (Example 2.6), but instead of using Adams and Kindred’s idea that basic state surfaces “generate” the geography of alternating knots, we use instead the constraint on the geography coming from essential surfaces described in Proposition 2.10. That is, for pretzel knots, essential surfaces play the role that basic state surfaces play for alternating knots. This means that we need to understand all possible essential surfaces in the complement of a pretzel knot.
@@ 6.1. Essential Surfaces and Edgepath Systems
Figure 21. Spanning surfaces corresponding to several edgepath systems: (a) the Seifert surface $\Lambda^{II}_{0}$, (b) the surface $\Lambda^{III}_{0}$, (c) the all-$A$ surface $\Lambda^{III}_{3}$, and (d) the all-$B$ surface $\Lambda^{III}_{4}$.
@@ 6.2. Computations on Edgepath Systems
The calculations above show that the Type I surface has $\Gamma^{-}(\Lambda^{I})=\frac{n+2}{2}-\frac{4}{n+1}$. For odd $n\geq 3$, we have $\Gamma^{-}(\Lambda^{I})\geq\frac{3}{2}$. Thus, any spanning surface that boundary compresses to the surface corresponding to $\Lambda^{I}$ must have $\Gamma^{-}(F)\geq 2$. Further, Table 1 shows that $\Gamma^{-}$ is at least $2$ for each of the Type II and Type III surfaces. Thus, Corollary 2.12 shows that $\hat{\gamma}_{3}^{-}(P_{n})\geq 2$. On the other hand, the all-$B$ surface corresponding to $\Lambda^{III}_{3}$ yields $\hat{\gamma}_{3}^{+}(P_{n})=0$. ∎
@@
Figure 22. The 33-dimensional geography of the knot $P_{3}$ is generated by the all-$A$ state surface at $(12,6)$ and the positively and negatively twisted Seifert surface at $(\pm 2,3)$. The Type I surface also lies at (2,3)(-2,3).
@@
Hatcher and Oertel’s algorithm goes further than just listing candidates the essential surfaces in the exterior of a Montesinos knot — it can actually identify which of these are essential. We chose not to use this part of Hatcher and Oertel’s work since the combinatorics necessary to pick out the essential surfaces from the candidates were more involved than simply computing the normal Euler and Betti numbers for all of the candidates. Nevertheless, computations using Dunfield’s implementation of Hatcher and Oertel’s algorithm [13] indicates that the essential surfaces for $P_{n}$ correspond to the edgepath systems $\Lambda^{I}$, $\Lambda^{II}_{0}$ (the Seifert surface), $\Lambda^{III}_{0}$, and $\Lambda^{III}_{3}$ (the all-$B$ surface).
7. Euler-Normalized Genus Gaps: Connected Sums
We conclude by investigating the behavior of the Euler-normalized non-orientable genus under connected sum. In particular, we prove the additivity of the Euler-normalized non-orientable 3-genus. The proof of Theorem 1.9 will follow as an immediate corollary.
@@ 7. Euler-Normalized Genus Gaps: Connected Sums
The Euler-normalized non-orientable $3$-genera $\hat{\gamma}^{\pm}_{3}$ are each additive under connected sum, i.e.
$\hat{\gamma}^{\pm}_{3}(K_{0}\#K_{1})=\hat{\gamma}^{\pm}_{3}(K_{0})+\hat{\gamma}^{\pm}_{3}(K_{1}).$
The proof of the lemma is, in essence, the same as the proof of subadditivity of the non-orientable $3$-genus in [9, Theorem 2.8] with two important differences:
@@
We write the proof in the positive case; the negative case is analogous. Let $F$ be a non-orientable spanning surface for $K_{0}\#K_{1}$ that realizes $\hat{\gamma}^{+}_{3}(K_{0}\#K_{1})$.
Using standard arguments (see, for example, [14, §5]), we may decompose $F$ as $F_{0}\natural F_{1}$, where $F_{i}$ is a spanning surface for $K_{i}$. If $F$ realizes $\hat{\gamma}^{+}_{3}(K_{0}\#K_{1})$, then so does $\operatorname{tw}^{k}_{+}(F)$, the surface obtained from the addition of $k$ positive twisted bands. Thus, we may assume that the surfaces $F_{i}$ are non-orientable. It suffices to prove that, for $i=0,1$, we have $\hat{\gamma}^{+}_{3}(K_{i})=\Gamma^{+}(F_{i})$.
@@
If $e(F_{0}^{*})\leq e(F_{0})$, then add $k=e(F_{0})-e(F_{0}^{*})$ positive twisted bands to $F_{0}^{*}$, yielding a surface $\operatorname{tw}^{k}_{+}(F_{0}^{*})$ with $e(\operatorname{tw}^{k}_{+}(F_{0}^{*}))=e(F_{0})$. As adding twisted bands does not change $\Gamma^{+}$, Equation (7.1) shows that $b_{1}(\operatorname{tw}^{k}_{+}(F_{0}^{*}))<b_{1}(F_{0})$. It follows that $b_{1}(\operatorname{tw}^{k}_{+}(F_{0}^{*})\natural F_{1})<b_{1}(F)$. Since $\operatorname{tw}^{k}_{+}(F_{0}^{*})\natural F_{1}$ and $F$ have the same normal Euler number, we see that $\Gamma^{+}(\operatorname{tw}^{k}_{+}(F_{0}^{*})\natural F_{1})<\Gamma^{+}(F)$. This contradicts the assumption that $F$ realizes $\hat{\gamma}^{+}_{3}(K_{0}\#K_{1})$.
@@
As in Section 6, consider the pretzel knot P=P(3,3,3)P=P(-3,3,3). We know that $\hat{\gamma}^{-}_{3}(P)=2$, while $\hat{\gamma}^{-}_{4}(P)=0$ since $P$ is slice. Let $K_{n}=\#_{n}P$. On one hand, Lemma 7.1 shows that $\hat{\gamma}^{-}_{3}(K_{n})=2n$. On the other, the knot $K_{n}$ is slice, so $\hat{\gamma}^{-}_{4}(K_{n})=0$. The theorem follows. ∎

### v2

@@ 5. Ordinary Genus Gaps for Pretzel Knots
Proof of Proposition 1.6.
The fact that $\gamma_{3}(P)=2n$ follows directly from the work of Ichihara and Mizushima [25, Theorem 1.2].
@@
As mentioned at the beginning of this section, despite the large gaps between $\gamma_{3}$ and $\gamma_{4}$ for the knots in Proposition 1.6, these examples do not exhibit similarly large gaps between $d^{\pm}_{3}$ and $d^{\pm}_{4}$.
@@
$d^{\pm}_{3}(P)-d^{\pm}_{4}(P)\leq 2.$
We conjecture that $d^{\pm}_{3}(P)-d^{\pm}_{4}(P)=0$. The conjecture would hold if the Möbius band filling constructed in the proof of Proposition 1.6 were the unique such band.
@@
Let $F_{A}$ (resp. $F_{B}$) be the surfaces for $P$ depicted in Figure 12. Direct computations on the diagrams yield
$e(F_{A})=2k+\sum_{i=1}^{n-1}p_{i}\quad\text{and}\quad e(F_{B})=-4r-2-\sum_{i=1}^{n-1}p_{i},$
as well as
@@
Figure 12. The surfaces $F_{A}$ (left) and $F_{B}$ (right) used in the proof of Proposition 5.1.
@@
6. Spanning Surface Defect Gaps for Slice Pretzel Knots
In this section, we prove Proposition 1.7 by exhibiting an infinite family of non-alternating slice pretzel knots with negative Euler-normalized non-orientable $3$-genus 22. The techniques deployed in the proof are interesting more for their suggestion of what work would be necessary for the computation of the spanning surface defect of adequate knots and the proof of Conjecture 1.4 than for the result itself which, as noted above, would also follow from Theorems 1.2 and 1.3. The idea is reminiscent of the computation of the $3$-dimensional geography of the trefoil, but instead of using Adams and Kindred’s idea that basic state surfaces “generate” the geography of alternating knots, we use instead the constraint on the geography coming from essential surfaces described in Proposition 2.2. That is, for pretzel knots, essential surfaces play the role that basic state surfaces play for alternating knots. This means that we need to understand all possible essential surfaces in the complement of a pretzel knot.
@@ 6.1. Essential Surfaces and Edgepath Systems
Figure 17. Spanning surfaces corresponding to several edgepath systems: (a) the Seifert surface $\Lambda^{II}_{0}$, (b) the surface $\Lambda^{III}_{0}$, (c) the surface $\Lambda^{III}_{3}$, and (d) the surface $\Lambda^{III}_{4}$.
@@ 6.2. Computations on Edgepath Systems
The calculations above show that the Type I surface has $\Gamma^{-}(\Lambda^{I})=\frac{n+2}{2}-\frac{4}{n+1}$. For odd $n\geq 3$, we have $\Gamma^{-}(\Lambda^{I})\geq\frac{3}{2}$. Thus, any spanning surface that boundary compresses to the surface corresponding to $\Lambda^{I}$ must have $\Gamma^{-}(F)\geq 2$. Further, Table 1 shows that $\Gamma^{-}$ is at least $2$ for each of the Type II and Type III surfaces. Thus, Corollary 2.4 shows that $d_{3}^{-}(P_{n})\geq 2$. On the other hand, the surface corresponding to $\Lambda^{III}_{3}$ yields $d_{3}^{+}(P_{n})=0$. ∎
@@
Figure 18. The $3$-dimensional geography of the knot $P_{3}$ is generated by surface at $(12,6)$ and the positively and negatively twisted Seifert surface at $(\pm 2,3)$. The Type I surface also lies at (2,3)(-2,3).
@@
Hatcher and Oertel’s algorithm goes further than just listing candidates the essential surfaces in the exterior of a Montesinos knot — it can actually identify which of these are essential. We chose not to use this part of Hatcher and Oertel’s work since the combinatorics necessary to pick out the essential surfaces from the candidates were more involved than simply computing the normal Euler and Betti numbers for all of the candidates. Nevertheless, computations using Dunfield’s implementation of Hatcher and Oertel’s algorithm [11] indicates that the essential surfaces for $P_{n}$ correspond to the edgepath systems $\Lambda^{I}$, $\Lambda^{II}_{0}$ (the Seifert surface), $\Lambda^{III}_{0}$, and $\Lambda^{III}_{3}$.
7. Spanning Surface Defect Gaps via Connected Sums
We conclude by investigating the behavior of the $3$-dimensional spanning surface defect under connected sum. The proof of Theorem 1.8 will follow as an immediate corollary.
@@ 7. Spanning Surface Defect Gaps via Connected Sums
The $3$-dimensional spanning surface defects $d^{\pm}_{3}$ are each additive under connected sum, i.e.
$d^{\pm}_{3}(K_{0}\#K_{1})=d^{\pm}_{3}(K_{0})+d^{\pm}_{3}(K_{1}).$
The proof of the lemma is, in essence, the same as the proof of subadditivity of the non-orientable $3$-genus in [8, Theorem 2.8] with two important differences:
@@
We write the proof in the positive case; the negative case is analogous. Let $F$ be a non-orientable spanning surface for $K_{0}\#K_{1}$ that realizes $d^{+}_{3}(K_{0}\#K_{1})$.
Using standard arguments (see, for example, [12, §5]), we may decompose $F$ as $F_{0}\natural F_{1}$, where $F_{i}$ is a spanning surface for $K_{i}$. If $F$ realizes $d^{+}_{3}(K_{0}\#K_{1})$, then so does $\operatorname{tw}^{k}_{+}(F)$, the surface obtained from the addition of $k$ positive twisted bands. Thus, we may assume that the surfaces $F_{i}$ are non-orientable. It suffices to prove that, for $i=0,1$, we have $d^{+}_{3}(K_{i})=\Gamma^{+}(F_{i})$.
@@
If $e(F_{0}^{*})\leq e(F_{0})$, then add $k=e(F_{0})-e(F_{0}^{*})$ positive twisted bands to $F_{0}^{*}$, yielding a surface $\operatorname{tw}^{k}_{+}(F_{0}^{*})$ with $e(\operatorname{tw}^{k}_{+}(F_{0}^{*}))=e(F_{0})$. As adding twisted bands does not change $\Gamma^{+}$, Equation (7.1) shows that $b_{1}(\operatorname{tw}^{k}_{+}(F_{0}^{*}))<b_{1}(F_{0})$. It follows that $b_{1}(\operatorname{tw}^{k}_{+}(F_{0}^{*})\natural F_{1})<b_{1}(F)$. Since $\operatorname{tw}^{k}_{+}(F_{0}^{*})\natural F_{1}$ and $F$ have the same normal Euler number, we see that $\Gamma^{+}(\operatorname{tw}^{k}_{+}(F_{0}^{*})\natural F_{1})<\Gamma^{+}(F)$. This contradicts the assumption that $F$ realizes $d^{+}_{3}(K_{0}\#K_{1})$.
@@
As in Section 6, consider the pretzel knot P=P(3,3,3)P=P(-3,3,3). We know that $d^{-}_{3}(P)=2$, while $d^{-}_{4}(P)=0$ since $P$ is slice. Let $K_{n}=\#_{n}P$. On one hand, Lemma 7.1 shows that $d^{-}_{3}(K_{n})=2n$. On the other, the knot $K_{n}$ is slice, so $d^{-}_{4}(K_{n})=0$. The theorem follows. ∎